# Speech Bandwidth Expansion Via High Fidelity Generative Adversarial Networks

### v1

@@ 2 Related Work
In [27], to to model the distribution of the target high-resolution signal conditioned on the log-scale mel-spectrogram of the low-resolution signal researchers utilized the WaveNet [28] model. Furthermore, [29] studied the use of the Wave-U-Net architecture for speech enhancement.
@@
A diffusion-based generative model designed to perform robust audio super-resolution across a wide range of audio types named AudioSR [35]. AudioSR designed specifically to enhance the quality of sound effects, music, and speech.
@@ 3 Methodology
Given a dataset $\mathcal{D}=\{(x_{1},\hat{x}_{1}),(x_{2},\hat{x}_{2}),\ldots,(x_{M},\hat{x}_{M})\}$, where each pair consists of the same speech signal sampled at different frequencies, we define $\hat{x}_{m}$ as the signal sampled at a lower frequency $F_{low}$ with bandwidth $B_{low}$, and $x_{m}$ as the signal sampled at a higher frequency $F_{high}$ with bandwidth $B_{high}$. Both signals originate from the same source. Mathematically, $F_{high}=\mathbf{s}\times F_{low}$ and $B_{high}=\mathbf{s}\times B_{low}$, where $\mathbf{s}$ is the super-resolution/upsampling ratio. Formally, we represent $x_{m}$ as $x_{m}=\{x_{1},x_{2},\ldots,x_{N}\}$ and $\hat{x}_{m}$ as $\hat{x}_{m}=\{\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{\lfloor\frac{N}{\mathbf{s}}\rfloor}\}$.
@@ 4.3 Experimental Setup
All the models were trained on mel-spectrogram input, calculated using $80$ mel-filter bank with an FFT size of $1024$, a window size of $1024$, and a hop length of $256$. For the generator loss terms, we set $\lambda_{adv}=1.1$, $\lambda_{mel}=50$, and $\lambda_{feat}=2$. In addition to that, the AdamW optimizer was used with $\beta_{1}=0.8$ and $\beta_{2}=0.999$, with a weight decay $\lambda=0.01$, and the learning rate used on each epoch calculated using the equation 9, where $lr_{init}$ is the initial learning rate set to $1.5\times 10^{-4}$, $epoch$ is the current training epoch, and $\gamma$ is the learning rate decaying factor set to $0.999$.

### v2

@@ 2 Related Work
In [27], to model the distribution of the target high-resolution signal conditioned on the log-scale mel-spectrogram of the low-resolution signal researchers utilized the WaveNet [28] model. Furthermore, [29] studied the use of the Wave-U-Net architecture for speech enhancement.
@@
A diffusion-based generative model designed to perform robust audio super-resolution across a wide range of audio types named AudioSR proposed in [35]. AudioSR designed specifically to enhance the quality of sound effects, music, and speech.
@@ 3 Methodology
Given a dataset $\mathcal{D}=\{(x_{1},\hat{x}_{1}),(x_{2},\hat{x}_{2}),\ldots,(x_{M},\hat{x}_{M})\}$, where each pair consists of the same speech signal sampled at different frequencies, we define $\hat{x}_{m}$ as the signal sampled at a lower frequency $F_{low}$ with bandwidth $B_{low}$, and $x_{m}$ as the signal sampled at a higher frequency $F_{high}$ with bandwidth $B_{high}$. Both signals originate from the same source. Mathematically, $F_{high}=\mathbf{s}\times F_{low}$ and $B_{high}=\mathbf{s}\times B_{low}$, such that $\mathbf{s}>1$, where $\mathbf{s}$ is the super-resolution/upsampling ratio. Formally, we represent $x_{m}$ as $x_{m}=\{x_{1},x_{2},\ldots,x_{N}\}$ and $\hat{x}_{m}$ as $\hat{x}_{m}=\{\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{\lfloor\frac{N}{\mathbf{s}}\rfloor}\}$.
@@ 4.3 Experimental Setup
All the models were trained on mel-spectrogram input, calculated using $80$ mel-filter bank with an FFT size of $1024$, a window size of $1024$, and a hop length of $256$. For the generator loss terms, we set $\lambda_{adv}=1.1$, $\lambda_{mel}=50$, and $\lambda_{feat}=2$. In addition to that, the AdamW optimizer was used with $\beta_{1}=0.8$ and $\beta_{2}=0.999$, with a weight decay $\lambda=0.01$, and the learning rate used on each epoch calculated using Equation 9, where $lr_{init}$ is the initial learning rate set to $1.5\times 10^{-4}$, $epoch$ is the current training epoch, and $\gamma$ is the learning rate decaying factor set to $0.999$.